Live GMAT classes with Target Test Prep start Nov 21 and Nov 22. First class free.

See classes
Target Test Prep · GMAT

Choose how you want to prepare

Learn live with an expert or move at your own pace. Every option includes the complete TTP study system.

Newsweek Readers’ Choice Awards 2026 winner: Target Test Prep, #1 Test Prep

Target Test Prep was voted #1 Best Test Prep in the 2026 Newsweek Readers’ Choice™ Awards.

★★★★★5.0563 reviews
GMATLiveTeachStarts Nov 22
Josh Braslow, Target Test Prep GMAT expert
LIVE ONLINE CLASSES

Get Ready for GMAT Test Day Faster with Live Online Classes

with Josh Braslow, 20+ years in test prep

Nov 22 · Josh Braslow
Sun · 9:30 AM to 12:30 PM ET
Included
40 hours of live online classes + 6 months of TTP OnDemand
  • Attend the first class for free
  • Every class is recorded, so you never fall behind
View classes & enroll
Limited seats availableTarget Test Prep
EALiveTeachOnDemandStart anytime
EXECUTIVE ASSESSMENT

Target Test Prep EA OnDemand

Self-paced EA prep. Study on your schedule.

EXECUTIVE ASSESSMENT

with

165+ EA score guarantee
$05-day trial no automatic billing
Schedule
Included
40 hours of live online classes plus six months of access to the complete TTP EA OnDemand course.
  • 165+ EA Score Guarantee
  • 4,100+ Quant, Verbal, and Integrated Reasoning practice questions
  • 400+ hours of in-depth video lessons
  • 3,000+ step-by-step video solutions
View EA class & enroll Start free 5-day trial
Limited cohort · enrollment openTrial includes full course accessTarget Test Prep
GMATOnDemandStart anytime
SELF-PACED MASTERCLASS

Target Test Prep GMAT OnDemand

Complete access from day one. Study on your schedule.

715+ score guarantee
$0to start then $127/mo
  • Personalized study plan and analytics
  • Thousands of lessons and practice questions

important gmat probability problem plz help

Expert replies
by quantskillsgmat » Tue Mar 06, 2012 11:09 pm
Q baseball world series matches 2 teams against each other in best of seven series.The first team to win four games wins the series and no subsequent games are played.If you have no special information about either of two teams,what is probability that world series will consist of fewer than 7 games.
a)0.125 b)0.25 c)0.3125 d)0.6875 e)0.75
Join the discussion
Source: — Problem Solving |

by kul512 » Wed Mar 07, 2012 4:46 am
As we don't have any specific information about a team, probability of wining for any team=1/2

We will calculate the probability of winning of team A. Then total result will be multiplied by 2, because both team have equal probability of winning the match. Minimum 4 matches are required to win the series and according to question we will get maximum of 6 matches.

4 matches-
Probability of team A winning the series in 4 matches(it wins all 4 matches)= (1/2)^4=(1/16)

5 matches-
5 matches will be required only if team A doesn't win 4 consecutive matches.
this is possible in following ways-
WWWLW
WWLWW
WLWWW
LWWWW
total=4

We can use permutation formula of arranging the 5 things in which 4 are identical-
5!/4!=5, but this also includes the case WWWWL, in that case match will be finished in 4 matches only hence we will deduct this one case (when series will be over in 4 matches.), so total matches =5-1=4

so probability= 4*(1/2)^5=1/8

6 matches
Last case will be when 6 matches are required, so total 4 W and 2 L, Total possible arrangement=6!/4!2!=15
But in Five cases (as mentioned above) series will be over in less than 6 matches. So total cases when series will be over in exactly 6 matches= 15-5=10

So probability= 10*(1/2)^6=5/32
total probability= 1/16 + 1/8 +5/32 = 11/32

Same all cases will be true for team B also, so total probability= 2*(11/32)=11/16= 0.6875

[spoiler]Ans (d)[/spoiler]
Sometimes there is very fine line between right and wrong: perspective.
Join the discussion

by GMATGuruNY » Wed Mar 07, 2012 4:56 am
Private tutor exclusively for the GMAT and GRE, with over 20 years of experience.
Followed here and elsewhere by over 1900 test-takers.
I have worked with students based in the US, Australia, Taiwan, China, Tajikistan, Kuwait, Saudi Arabia -- a long list of countries.
My students have been admitted to HBS, CBS, Tuck, Yale, Stern, Fuqua -- a long list of top programs.

As a tutor, I don't simply teach you how I would approach problems.
I unlock the best way for YOU to solve problems.

For more information, please email me (Mitch Hunt) at [email protected].
Student Review #1
Student Review #2
Student Review #3
Join the discussion

by krusta80 » Wed Mar 07, 2012 7:14 am
quantskillsgmat wrote:Q baseball world series matches 2 teams against each other in best of seven series.The first team to win four games wins the series and no subsequent games are played.If you have no special information about either of two teams,what is probability that world series will consist of fewer than 7 games.
a)0.125 b)0.25 c)0.3125 d)0.6875 e)0.75
STEP 1: What are we counting? We are counting distinct world-series game arrangements.

STEP 2: What defines a world-series game arrangmeent?

A world-series game arrangement is a specific ordering of world-series games. Each game has two possible values: {A,B}, where A represents one team winning, and B represents the other team winning. The number of series games varies from 4 to 7 games, since the series ends when we find either 4 A's or 4 B's.

STEP 3: Let's count the total number of game arrangements.

Since either team can win (or lose) the series the same number of ways, we can take advantage of the available symmetry and only look at possibilities that lead to team A winning the series and then mulitply it by 2.

Since we have the restriction that A must win the last game, the rest of the series games will have exactly three A's from which to create our groups of games. The number of B's to choose from will equal the series length minus four.

Instead of calculating the probability for each series length, we can simply calculate the probability of the series going seven games, and then subtracting that from 1.

p(A winning in 7)
------A

How many words of six ordered letters can we have with three A's and three B's in it? The answer is 6!/3!/3! = 20

Each distinct order has a probability of 1/2^7 = 1/128 of occuring, so our total probability of A winning in seven games is 20/128 = 5/32, which means there is a 10/32 probability of either A or B winning in seven games.

p(less than 7) = 1 - p(seven) = 22/32 = 11/16

D
Join the discussion